You are analyzing two assets: collectible LEGO sets, and stock of Apple. In the last 5 years, LEGOs have had an annual volatility of 5%, annual return of 6%, and a CAPM beta (the correlation coefficient between the asset and the market risk-premium) of 1.6. Apple has had an annual volatility of 10%, an annual return of 8%, and a CAPM beta of 1.2.

1) If the risk-premium of the market is currently 7% and the risk-free rate is 2%, what is the expected return of LEGO sets this year in %?

(HINT: remember that CAPM beta is the sensitivity to the market risk premium: To calculate expected return, use E(r) = rf + beta* (market – rf) as opposed to E(r) = beta*market)
Answer #1

Expected return = rf +beta*(required return of market-risk free rate)

risk free rate= 2%

Risk premium of the market = Required returns of the market - riskfree rate

7% = rM - 2%

rM = 5%

Expected return = 0.02 +1.6(0.05-0.02) = 6.8%